# Newton’s method for nonlinear mappings into vector bundles

### v1

@@ 2. Preliminaries
The concept of a connection of a vector bundle $p:\mathcal{E}\to\mathcal{M}$ is fundamental to differential geometry. It is an additional geometric structure imposed on $\mathcal{E}$ that describes in an infinitesimal way, how neighboring fibres are related. The most prominent example is the Levi-Civita connection on $\mathcal{E}=T\mathcal{X}$ (see e.g. [13, VIII §4]) for a Riemannian manifold $\mathcal{X}$. Connections give rise to further concepts like the covariant derivative, curvature, geodesics, and parallel transports. They are described in the literature in various equivalent ways (cf. e.g. [13, IV §3 or X §4 or XIII]). For the purpose of our paper we choose a formulation that emphasizes the idea of a connection map $Q_{e}:T_{e}\mathcal{E}\to E_{p(e)}$, which then induces a splitting of $T_{e}\mathcal{E}$ into a vertical and a horizontal subspace.
It is well known that the kernel of $p^{\prime}:T\mathcal{E}\rightarrow T\mathcal{M}$ canonically defines the vertical subbundle $V\mathcal{E}$ of $\mathcal{E}$ [13, IV §3]. Its fibres are closed linear subspaces, called the vertical subspaces $V_{e}:=\ker p^{\prime}(e)\subset T_{e}\mathcal{E}$. We can identify $E_{p(e)}\cong V_{e}$ canonically by the isomorphism $\operatorname{\kappa}_{e}:E_{p(e)}\rightarrow V_{e}$, given by $w\mapsto\frac{d}{dt}(e+tw)|_{t=0}$. Elements of $V_{e}$ are represented in trivializations by pairs of the form $(0_{y},\delta e_{\tau})\in T_{y}\mathcal{M}\times\mathbb{E}$.
Using this identification, we define a connection map $Q:T\mathcal{E}\rightarrow\mathcal{E}$ as a fibrewise projection onto $\mathcal{E}$. More precisely, $Q\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}\mathcal{E}))$, i.e., for each $e\in\mathcal{E}$ we get a continuous linear map $Q_{e}\in L(T_{e}\mathcal{E},E_{p(e)})$, such that $Q_{e}\operatorname{\kappa}_{e}=Id_{E_{p(e)}}$. The kernels $H_{e}:=\ker Q_{e}$ are called horizontal subspaces, collected in the horizontal subbundle $H\mathcal{E}$. While $V\mathcal{E}$ is given canonically, a choice of $Q$ (or equivalently $H\mathcal{E}$), imposes additional geometric structure on $\mathcal{E}$.
Using the representations $(y,e_{\tau})\in\mathcal{M}\times\mathbb{E}$ for $e\in\mathcal{E}$ and $(\delta y,\delta e_{\tau})\in T_{x}\mathcal{M}\times\mathbb{E}$ for $\delta e\in T_{e}\mathcal{E}$, the representation for a linear connection map $Q$ in trivializations can be written in the form
@@
where $B_{y,\tau}:\mathbb{E}\to L(T_{y}\mathcal{M},\mathbb{E})$ assigns a linear mapping $B_{y,\tau}(e_{\tau})$ to each $e_{\tau}\in\mathcal{E}$. Since $Q_{e,\tau}(0_{y},\delta e_{\tau})=\delta e_{\tau}$, we see that $Q_{e,\tau}$ indeed represents a projection onto $V_{y}$. To reflect the linearity of the fibres we want the connection to be linear (fibrewise with respect to $e$). To this end, we consider the fibrewise scaling $m_{s}:\mathcal{E}\rightarrow\mathcal{E},e\mapsto se$ and require the condition $Q\circ m_{s}^{\prime}=m_{s}\circ Q$ for all $s\in\mathbb{K}$, which reads in trivializations:
$Q_{se,\tau}(\delta y,s\delta e_{\tau})=sQ_{e,\tau}(\delta y,\delta e_{\tau})\quad\forall s\in\mathbb{K}.$
It can be shown that $Q$ is linear, if and only if the mapping $e_{\tau}\to B_{y,\tau}(e_{\tau})$ is linear, or put differently, the mapping $(\delta y,e_{\tau})\to B_{y,\tau}(e_{\tau})\delta y$ is bilinear. Alternatively, for a vector bundle chart $\Phi$ we have a representation $B_{y,\Phi}:\mathbb{M}\times\mathbb{E}\to\mathbb{E}$.
If $\mathcal{E}=T\mathcal{X}$ in classical Riemannian geometry, where $Q$ is given by the Levi-Civita connection, the bilinear mapping $B_{x,\Phi}:\mathbb{X}\times\mathbb{X}\to\mathbb{X}$ is represented by Christoffel symbols, and it is symmetric, i.e. $B_{x,\Phi}(v_{i})w_{i}=B_{x,\Phi}(w_{i})v_{i}$, in natural tangent bundle charts.
@@
If $V_{y}^{-1}$ is defined via pointwise inverses of a vector transport $V_{y}\in\Gamma(\mathcal{M}\times E_{y},\mathcal{E})$, then
@@
By differentiation of ${\langle{V^{-1}_{y}}\rangle}:\mathcal{E}\to E_{y}$ we obtain that ${\langle{V^{-1}_{y}}\rangle}^{\prime}(e):T_{e}\mathcal{E}\to E_{y}$ is a continuous linear map. Using (2) and $V_{y}^{-1}(y)=Id_{E_{y}}$, the representation of ${\langle{V_{y}^{-1}}\rangle}^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{E},E_{y}))$ in trivializations, where $V_{y,\tau}^{-1}:\mathcal{M}\to L(\mathbb{E},\mathbb{E})$, reads
${\langle{V_{y}^{-1}}\rangle}_{\tau}^{\prime}(e)(\delta y,\delta e_{\tau})=\delta e_{\tau}+((V_{y,\tau}^{-1})^{\prime}(y)\delta y)e_{\tau}\quad\forall(\delta y,\delta e_{\tau}),$
which is of the form (3), implying (4) and linearity. Let $\hat{e}\in E_{y}$. Since $\kappa_{e}(\hat{e})\in V_{e}\cong E_{y}$ we get the representation $\kappa_{e}(\hat{e})_{\tau}\sim(0_{y},\widehat{\delta e}_{\tau})$. Thus, we obtain $Q_{e}\kappa_{e}=Id_{E_{y}}$.
@@
Let $e\in\mathcal{E}$ and $y=p(e)$. We call a vector back-transport $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ consistent with a connection map $Q:T\mathcal{E}\to\mathcal{E}$ at $e$, if $Q_{e}={\langle{V^{-1}_{y}}\rangle}^{\prime}(e)$.
@@ 3. Definition of Newton’s method
Let $F:\mathcal{X}\rightarrow\mathcal{E}$ be a differentiable mapping between a Banach manifold $\mathcal{X}$ and a vector bundle $p:\mathcal{E}\to\mathcal{M}$. By the composition
$y:=p\circ F:\mathcal{X}\to\mathcal{M},\;y(x):=p(F(x))\in\mathcal{M},$
we can compute the base point $y(x)\in\mathcal{M}$ of $F(x)\in\mathcal{E}$ for given $x\in\mathcal{X}$. Consider the following root finding problem
@@
In contrast to the classical case, the linear space $E_{y(x)}$ in which $F(x)$ is evaluated now depends on $x$. Hence, iterative methods will have to deal with the situation that $E_{y(x)}$ changes during the iteration.
To derive Newton’s method for this problem we need to define a suitable Newton direction $\delta x\in T_{x}\mathcal{X}$. The tangent map of $F$ is a mapping $F^{\prime}:T\mathcal{X}\rightarrow T\mathcal{E}$. Since $F^{\prime}(x)\delta x\in T_{F(x)}\mathcal{E}$ and $F(x)\in E_{y(x)}$, these two quantities cannot be added, and thus the classical Newton equation $F^{\prime}(x)\delta x+F(x)=0$ is not well defined.
A connection map $Q:T\mathcal{E}\rightarrow\mathcal{E}$ on $\mathcal{E}$ allows us to resolve this issue. With its help we can define the following linear mapping at $x\in\mathcal{X}$, which maps into the correct space:
$Q_{F(x)}\circ F^{\prime}(x):T_{x}\mathcal{X}\to E_{y(x)}.$
@@
If $Q_{F(x)}\circ F^{\prime}(x)$ is invertible, then the Newton direction $\delta x\in T_{x}\mathcal{X}$ is given as the unique solution of this equation, and $\delta x=0$ if and only if $F(x)=0$.
The classical additive Newton update $x_{+}=x+\delta x$ is also not well defined since $x\in\mathcal{X}$ and $\delta x\in T_{x}\mathcal{X}$ cannot be added. To obtain a new iterate, we have to map $\delta x\in T_{x}\mathcal{X}$ back to the manifold $\mathcal{X}$, which can be done by a retraction, a popular concept, used widely in numerical methods on manifolds [1, Chap. 4]:
@@
A retraction on a manifold $\mathcal{X}$ is a smooth mapping $R:T\mathcal{X}\rightarrow\mathcal{X}$, where $R_{x}:T_{x}\mathcal{X}\to\mathcal{X}$ for a fixed $x\in\mathcal{X}$ satisfies the following properties:
@@
Thus, after successfully computing the Newton direction we use a retraction to generate the Newton step
@@ 4. Newton-Differentiability and Strict Differentiability
For the analysis of Newton’s method we will carry over two differentiability concepts, which are known on linear spaces, but slightly non-standard in the context of manifolds. The first is the so called Newton-differentiability [17, 20, 25], a concept which is tailored for the analysis of local convergence of Newton’s method, even for semismooth problems, the second one is strict differentiability [18, 21], a notion that is stronger than Fréchet differentiability, but weaker than continuous differentiability. As usual, we define these concepts by looking at the problem in local charts and showing invariance with respect to changes of charts.
@@
Consider $f:\mathcal{X}\to\mathcal{Y}$ where $\mathcal{X}$ and $\mathcal{Y}$ are both $C^{1}$ Banach manifolds. Let $(U,\phi)$ be a chart of $\mathcal{X}$ and $(V,\psi)$ be a chart of $\mathcal{Y}$. For $(U,\phi)$ and $(V,\psi)$ denote by $\Phi$ and $\Psi$ the corresponding vector bundle charts of $T\mathcal{X}$ and $T\mathcal{Y}$ according to Remark 2.1. We define the pullback
$f_{\phi,\psi}:=\psi\circ f\circ\phi^{-1}:\phi(U)\to\mathbb{Y}.$
Using this chart representation, we can define Newton-differentiability of $f$ with respect to $f^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},f^{*}T\mathcal{Y}))$ (which need not be the classical derivative) in the following way:
@@
We call $f$ Newton-differentiable at a point $x_{0}\in\mathcal{X}$ with respect to $f^{\prime}$ if the chart representations of $f$ and $f^{\prime}$ satisfy
$\lim_{\xi\to\xi_{0}}\frac{\|f^{\prime}_{\Phi,\Psi}(\xi)(\xi-\xi_{0})-(f_{\phi,\psi}(\xi)-f_{\phi,\psi}(\xi_{0}))\|_{\mathbb{Y}}}{\|\xi-\xi_{0}\|_{\mathbb{X}}}=0,$
where $\xi_{0}=\phi(x_{0})$ and $\xi\in\phi(U)$. We call $f^{\prime}$ a Newton-derivative of $f$ at $x_{0}$.
Be aware that Newton-derivatives are not unique, that’s why this definition refers to a given pair $f,f^{\prime}$. As we will see in the following discussion, local Lipschitz continuity of $f$ and local boundedness of $f^{\prime}$ are required to render Newton-differentiability independent of charts.
Analogously, we can define strict differentiability of $f:\mathcal{X}\to\mathcal{Y}$ in the following way:
@@
Let $f:\mathcal{X}\to\mathcal{Y}$. $f$ is called strictly differentiable at $x_{0}\in\mathcal{X}$, if there is $f^{\prime}(x_{0})\in L(T_{x_{0}}\mathcal{X},T_{f(x_{0})}\mathcal{Y})$, such that the chart representations of $f$ and $f^{\prime}$ satisfy the following condition:
For every $\varepsilon>0$ there exists a neighborhood $U$ of $\xi_{0}=\phi(x_{0})$ such that for $\xi,\eta\in U$
$\|f^{\prime}_{\Phi,\Psi}(\xi_{0})(\xi-\eta)-(f_{\phi,\psi}(\xi)-f_{\phi,\psi}(\eta))\|_{\mathbb{Y}}<\varepsilon\|\xi-\eta\|_{\mathbb{X}}.$
The mapping $f^{\prime}(x_{0})$ is called strict derivative of $f$ at $x_{0}$.

### v2

@@ 2. Preliminaries
The concept of a connection on a vector bundle $p:\mathcal{E}\to\mathcal{Y}$ is fundamental to differential geometry. It is an additional geometric structure imposed on $\mathcal{E}$ that describes in an infinitesimal way, how neighboring fibres are related. The most prominent example is the Levi-Civita connection on $\mathcal{E}=T\mathcal{X}$ (see, e.g., [14, VIII §4]) for a Riemannian manifold $\mathcal{X}$. Connections give rise to further concepts like the covariant derivative, curvature, geodesics, and parallel transports. They are described in the literature in various equivalent ways (cf., e.g., [14, IV §3 or X §4 or XIII]). For the purpose of our paper we choose a formulation that emphasizes the idea of a connection map $Q_{e}:T_{e}\mathcal{E}\to E_{p(e)}$, which then induces a splitting of $T_{e}\mathcal{E}$ into a vertical and a horizontal subspace.
It is well known that the kernel of $p^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}T\mathcal{Y}))$ canonically defines the vertical subbundle $V\mathcal{E}$ of $\mathcal{E}$ [14, IV §3]. Its fibres are closed linear subspaces, called the vertical subspaces $\mathrm{Vert}_{e}:=\ker p^{\prime}(e)\subset T_{e}\mathcal{E}$. We can identify $E_{p(e)}\cong\mathrm{Vert}_{e}$ canonically by the isomorphism $\operatorname{\kappa}_{e}:E_{p(e)}\rightarrow\mathrm{Vert}_{e}$, given by $w\mapsto\frac{d}{dt}(e+tw)|_{t=0}$. Elements of $\mathrm{Vert}_{e}$ are represented in trivializations by pairs of the form $(0_{y},\delta e_{\tau})\in T_{y}\mathcal{Y}\times\mathbb{E}$.
We define a connection map as a section $Q\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}\mathcal{E}))$, i.e., for each $e\in\mathcal{E}$ we get a continuous linear map $Q_{e}\in L(T_{e}\mathcal{E},E_{p(e)})$, with the additional property that $Q_{e}\operatorname{\kappa}_{e}=Id_{E_{p(e)}}$. The kernels $\mathrm{Hor}_{e}:=\ker Q_{e}$ are called horizontal subspaces, collected in the horizontal subbundle $H\mathcal{E}$. While $V\mathcal{E}$ is given canonically, a choice of $Q$ (or equivalently $H\mathcal{E}$), imposes additional geometric structure on $\mathcal{E}$. Using the identification $\mathrm{Vert}_{e}\cong E_{p(e)}$, we can view $Q$ as a fibrewise “projection” onto $\mathcal{E}$, with $\langle Q\rangle:T\mathcal{E}\rightarrow\mathcal{E}$. In trivializations $(y,e_{\tau})\in\mathcal{Y}\times\mathbb{E}$ for $e\in\mathcal{E}$ and $(\delta y,\delta e_{\tau})\in T_{y}\mathcal{Y}\times\mathbb{E}$ for $\delta e\in T_{e}\mathcal{E}$, the representation for a connection map $Q$ can be written in the form
@@
where $B_{y,\tau}:\mathbb{E}\to L(T_{y}\mathcal{Y},\mathbb{E})$ assigns a linear mapping $B_{y,\tau}(e_{\tau})$ to each $e_{\tau}\in\mathbb{E}$. Since $Q_{e,\tau}(0_{y},\delta e_{\tau})=\delta e_{\tau}$, we see that $Q_{e,\tau}$ indeed represents a projection onto $\mathrm{Vert}_{e}$. To reflect the linearity of the fibres we want the connection to be linear (fibrewise with respect to $e$). To this end, we consider the fibrewise scaling $m_{s}:\mathcal{E}\rightarrow\mathcal{E},e\mapsto se$ and require the condition $Q\circ m_{s}^{\prime}=m_{s}\circ Q$ for all $s\in\mathbb{R}$, which reads in trivializations:
$Q_{se,\tau}(\delta y,s\delta e_{\tau})=sQ_{e,\tau}(\delta y,\delta e_{\tau})\quad\forall s\in\mathbb{R}.$
It can be shown that $Q$ is linear, if and only if the mapping $e_{\tau}\mapsto B_{y,\tau}(e_{\tau})$ is linear, or put differently, the mapping $(\delta y,e_{\tau})\mapsto B_{y,\tau}(e_{\tau})\delta y$ is bilinear. Alternatively, for a vector bundle chart $\Phi$ we have a representation $B_{y,\Phi}:\mathbb{Y}\times\mathbb{E}\to\mathbb{E}$.
If $\mathcal{E}=T\mathcal{X}$ in classical Riemannian geometry, where $Q$ is given by the Levi-Civita connection, the bilinear mapping $B_{x,\Phi}:\mathbb{X}\times\mathbb{X}\to\mathbb{X}$ is represented by Christoffel symbols, and it is symmetric, i.e., $B_{x,\Phi}(v_{i})w_{i}=B_{x,\Phi}(w_{i})v_{i}$, in natural tangent bundle charts.
@@
If $V_{y}^{-1}$ is defined via fibrewise inverses of a vector transport $V_{y}\in\Gamma(\mathcal{Y}\times E_{y},\mathcal{E})$, then
@@
By differentiation of ${\langle{V^{-1}_{y}}\rangle}:\mathcal{E}\to E_{y}$ we obtain that ${\langle{V^{-1}_{y}}\rangle}^{\prime}(e):T_{e}\mathcal{E}\to E_{y}$ is a continuous linear map. Using (2) and $V_{y}^{-1}(y)=Id_{E_{y}}$, the representation of ${\langle{V_{y}^{-1}}\rangle}^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{E},E_{y}))$ in trivializations, where $V_{y,\tau}^{-1}:\mathcal{Y}\to L(\mathbb{E},\mathbb{E})$, reads
${\langle{V_{y}^{-1}}\rangle}_{\tau}^{\prime}(y,e_{\tau})(\delta y,\delta e_{\tau})=\delta e_{\tau}+((V_{y,\tau}^{-1})^{\prime}(y)\delta y)e_{\tau}\quad\forall(\delta y,\delta e_{\tau}),$
which is of the form (3), implying (4) and linearity. Let $\hat{e}\in E_{y}$. Since $\kappa_{e}(\hat{e})\in\mathrm{Vert}_{e}\cong E_{y}$ we get the representation $\kappa_{e}(\hat{e})_{\tau}\sim(0_{y},\widehat{\delta e}_{\tau})$. Thus, we obtain $Q_{e}\kappa_{e}=Id_{E_{y}}$.
@@
Let $e\in\mathcal{E}$ and $y=p(e)$. We call a vector back-transport $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$ consistent with a connection map $Q\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}\mathcal{E}))$ at $e$, if $Q_{e}={\langle{V^{-1}_{y}}\rangle}^{\prime}(e)$.
Remark 2.5.
In classical Riemannian geometry, the Levi-Civita connection (see, e.g., [14]) on $T\mathcal{X}$ is often used. This connection induces the parallel transport along geodesics and the exponential map, which may serve algorithmically as a vector transport and as a retraction. However, from al algorithmic point of view, these classical mappings, which are defined via integration, can be computationally very expensive. Then, instead of starting with a connection and deriving a vector transport from it, one can take the opposite route: start with a computationally tractable vector back-transport $V^{-1}_{y}\in\Gamma(\mathcal{L}(\mathcal{E},E_{y}))$, and derive a consistent connection $Q$ by differentiation.
@@ 3. Definition of Newton’s method
Let $F:\mathcal{X}\rightarrow\mathcal{E}$ be a differentiable mapping between a Banach manifold $\mathcal{X}$ and a vector bundle $p:\mathcal{E}\to\mathcal{Y}$. By the composition
$y:=p\circ F:\mathcal{X}\to\mathcal{Y},\;y(x):=p(F(x))\in\mathcal{Y},$
we can compute the base point $y(x)\in\mathcal{Y}$ of $F(x)\in\mathcal{E}$ for given $x\in\mathcal{X}$. Consider the following root finding problem
@@
In contrast to classical root finding problems, the linear space $E_{y(x)}$ in which $F(x)$ is evaluated now depends on $x$. Hence, when constructing iterative methods we will have to take into account that $E_{y(x)}$ changes during the iteration.
To derive Newton’s method for this problem we need to define a suitable Newton direction $\delta x\in T_{x}\mathcal{X}$. The tangent map of $F$ is a mapping $F^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},F^{*}T\mathcal{E}))$. Since $F^{\prime}(x)\delta x\in T_{F(x)}\mathcal{E}$ and $F(x)\in E_{y(x)}$, these two quantities cannot be added, and thus the classical Newton equation $F^{\prime}(x)\delta x+F(x)=0$ is not well defined.
A connection map $Q\in\Gamma(\mathcal{L}(T\mathcal{E},p^{*}\mathcal{E}))$ allows us to resolve this issue. With its help we can define the following linear mapping at $x\in\mathcal{X}$, which maps into the correct space:
$Q_{F(x)}\circ F^{\prime}(x):T_{x}\mathcal{X}\to E_{y(x)}.$ (6)
@@
If $Q_{F(x)}\circ F^{\prime}(x)$ is invertible, then the Newton direction $\delta x\in T_{x}\mathcal{X}$ is given as the unique solution of this equation, and $\delta x=0$ holds if and only if $F(x)=0$.
The classical additive Newton update $x_{+}=x+\delta x$ is also not well defined since $x\in\mathcal{X}$ and $\delta x\in T_{x}\mathcal{X}$ cannot be added. To obtain a new iterate, we have to map $\delta x\in T_{x}\mathcal{X}$ back to the manifold $\mathcal{X}$, which can be done by a retraction, a popular concept, used widely in numerical methods on manifolds (cf., e.g., [1, Chap. 4]).
@@
A $C^{1}$-retraction on a manifold $\mathcal{X}$ is a $C^{1}$-mapping $R:T\mathcal{X}\rightarrow\mathcal{X}$, where $R_{x}:T_{x}\mathcal{X}\to\mathcal{X}$ for a fixed $x\in\mathcal{X}$ satisfies the following properties:
@@
Thus, after successfully computing the Newton direction we use a local retraction to generate the Newton step
@@ 4. Newton-Differentiability and Strict Differentiability
Classically, quadratic convergence of Newton’s method is analysed for continuously differentiable mappings, where the derivative satisfies a Lipschitz condition. However, in recent years, weaker smoothness concepts have become popular, which are sufficient to show local subperlinear convergence. For the analysis of Newton’s method we will discuss two of these differentiability concepts, which are known on linear spaces, but slightly non-standard in the context of manifolds. The first is the so called Newton-differentiability [18, 21, 27], a concept which is tailored for the analysis of local convergence of Newton’s method, even for semismooth problems, the second one is strict differentiability [19, 23], a notion that is stronger than Fréchet differentiability, but weaker than continuous differentiability, and we will use it to study the simplified Newton method, which is a building block of our globalilzation scheme, introduced below. As usual, we define these concepts by looking at the problem in local charts and showing invariance with respect to changes of charts.
@@
Consider $F:\mathcal{X}\to\mathcal{Y}$ where $\mathcal{X}$ and $\mathcal{Y}$ are both $C^{1}$-Banach manifolds. Let $(U,\phi)$ be a chart of $\mathcal{X}$ and $(V,\psi)$ be a chart of $\mathcal{Y}$. Consider the following representation of $F$ in charts
$F_{\psi,\phi}:=\psi\circ F\circ\phi^{-1}:\phi(U)\to\mathbb{Y}.$
Consider also a fibrewise linear mapping $F^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},F^{*}T\mathcal{Y}))$ with chart representations
$(F_{\psi,\phi})^{\prime}:\phi(U)\to L(\mathbb{X},\mathbb{Y}).$
As a first step, we will state our differentiability concepts for these chart representations. $F_{\psi,\phi}$ is called Newton-differentiable at $\xi_{0}\in\phi(U)$ with respect to a Newton-derivative $(F_{\psi,\phi})^{\prime}$ if
$\lim_{\xi\to\xi_{0}}\frac{\|(F_{\psi,\phi})^{\prime}(\xi)(\xi-\xi_{0})-(F_{\psi,\phi}(\xi)-F_{\psi,\phi}(\xi_{0}))\|_{\mathbb{Y}}}{\|\xi-\xi_{0}\|_{\mathbb{X}}}=0.$ (7)
Newton-derivatives are, in contrast to classical derivatives, not unique, since the linearization $(F_{\psi,\phi})^{\prime}(\xi)$ is evaluated at the moving point $\xi$. This is the reason, why our definition refers to a given pair $F_{\psi,\phi},\;(F_{\psi,\phi})^{\prime}$.
Example 4.1.
The real function $F(x)=|x|$ is Newton-differentiable at $x_{0}=0$ with respect to $F^{\prime}$, if there is a neighbourhood $U$ of $x_{0}$, such that $F^{\prime}(x)=\mathrm{sgn}\,x$ for all $x\in U\setminus\{x_{0}\}$.
We call $F_{\psi,\phi}$ strictly differentiable at a point $\xi_{0}\in\phi(U)$ if there is $(F_{\psi,\phi})^{\prime}(\xi_{0})\in L(\mathbb{X},\mathbb{Y})$, such that the following condition holds:
For every $\varepsilon>0$ there exists a neighborhood $U_{\varepsilon}$ of $\xi_{0}$ such that
$\|(F_{\psi,\phi})^{\prime}(\xi_{0})(\xi-\eta)-(F_{\psi,\phi}(\xi)-F_{\psi,\phi}(\eta))\|_{\mathbb{Y}}<\varepsilon\|\xi-\eta\|_{\mathbb{X}}\quad\forall\xi,\eta\in U_{\varepsilon}.$ (8)
Strict differentiability is a slightly stronger notion than Fréchet differentiability, but weaker than continuous differentiability. By the triangle inequality, we see that strictly differentiable mappings are locally Lipschitz continuous near $\xi_{0}$ with a Lipschitz constant that is close to $\|(F_{\psi,\phi})^{\prime}(\xi_{0})\|_{\mathbb{X}\to\mathbb{Y}}$. In linear spaces it is well known, and also not hard to verify, that continuous differentiability implies Newton-differentiability (w.r.t. the classical derivative), as well as strict differentiability.
Our second step is to show independence of these definitions of the choice of charts. To this end we need the following chain rules for Newton- and strict differentiability in linear spaces:
Lemma 4.2.
Let $\mathbb{X},\;\mathbb{Y}$ and $\mathbb{Z}$ be normed linear spaces. Consider $h:\mathbb{X}\to\mathbb{Y}$ and $g:\mathbb{Y}\to\mathbb{Z}$.
  1. (i)
    Let $h$ and $g$ be Newton-differentiable at $\xi_{0}$ and $h(\xi_{0})$ with respect to Newton-derivatives $h^{\prime}$ and $g^{\prime}$ in open sets $V$ and $U$, respectively, where $\xi_{0}\in U\subset\mathbb{X}$, $h(U)\subset V\subset\mathbb{Y}$. Assume that $h$ is Lipschitz continuous in $U$ and $g^{\prime}$ is locally bounded in $V$. Then $g\circ h:\mathbb{X}\to\mathbb{Z}$ is Newton-differentiable with respect to the Newton-derivative $g^{\prime}(h(\cdot))h^{\prime}(\cdot):U\to L(\mathbb{X},\mathbb{Z})$.
  2. (ii)
    Let $h$ and $g$ be strictly differentiable at $\xi_{0}$ and $h(\xi_{0})$ with derivatives $h^{\prime}(\xi_{0})$ and $g^{\prime}(h(\xi_{0}))$, respectively. Then $g\circ h:\mathbb{X}\to\mathbb{Z}$ is strictly differentiable with derivative $g^{\prime}(h(\xi_{0}))h^{\prime}(\xi_{0})\in L(\mathbb{X},\mathbb{Z})$.
Proof.
The proofs of both results are rather standard, but difficult to retrieve in the literature. For Newton-differentiabilty, a proof can be found in the lecture notes [11], for strict differentiability it is stated but not proved in [19]. Thus, for the sake of completeness we provide concise proofs here.
Let $u,v,w\in\mathbb{X}$. To cover Newton-differentiability (7), consider $u=v=\xi\in U$ and $w=\xi_{0}$, for strict differentiability (8) consider $u=\xi_{0}$, $v=\xi$ and $w=\eta$.
Let us split the remainder term of interest as follows:
$r(u,v,w)$ $:=g^{\prime}(h(u))h^{\prime}(u)(v-w)-(g(h(v))-g(h(w)))=r_{g}(u,v,w)+g^{\prime}(h(u))r_{h}(u,v,w)$ (9)
$r_{g}(u,v,w)$ $:=g^{\prime}(h(u))(h(v)-h(w))-(g(h(v))-g(h(w)))$
$r_{h}(u,v,w)$ $:=h^{\prime}(u)(v-w)-(h(v)-h(w)).$
Since $g^{\prime}$ is locally bounded by assumption in the Newton-differentiable case and $u=x_{0}$ is fixed in the strictly differentiable case, there exists a local constant $C>0$ such that $\|g^{\prime}(h(u))\|_{\mathbb{Y}\to\mathbb{Z}}\leq C$. Thus, we obtain in both cases
$\|r(u,v,w)\|_{\mathbb{Z}}\leq\|r_{g}(u,v,w)\|_{\mathbb{Z}}+C\|r_{h}(u,v,w)\|_{\mathbb{Z}}.$
By definition of Newton-differentiability/strict differentiability of $g$ and $h$ and local Lipschitz continuity of $h$ (which holds by assumption in the first case and as a consequence of strict differentiability in the second case) we obtain that for each $\varepsilon>0$ there is a neighborhood $U_{\varepsilon}$ of $\xi_{0}$, such that
$\|r_{g}(u,v,w)\|_{\mathbb{Z}}$ $\leq\varepsilon\|h(v)-h(w)\|_{\mathbb{Y}}\leq\varepsilon L_{h}\|v-w\|_{\mathbb{X}}$
$\|r_{h}(u,v,w)\|_{\mathbb{Z}}$ $\leq\varepsilon\|v-w\|_{\mathbb{X}},$
for all choices of $u,v,w\in U_{\varepsilon}$ corresponding to the respective differentiability concepts. This yields our desired results. ∎
Proposition 4.3.
Consider $F:\mathcal{X}\to\mathcal{Y}$.
  • i)
    Let $F$ be locally Lipschitz continuous and $F^{\prime}\in\Gamma(\mathcal{L}(T\mathcal{X},F^{*}T\mathcal{Y}))$ be locally bounded. Then the criterion (7) for Newton-differentiability is independent of the choice of charts.
  • ii)
    The criterion (8) for strict differentiability of $F$ is independent of the choice of charts.
Proof.
$\psi_{j}\circ F\circ\phi_{j}^{-1}=t_{ji}^{\mathcal{Y}}\circ(\psi_{i}\circ F\circ\phi_{i}^{-1})\circ t_{ij}^{\mathcal{X}}.$
the changes of charts $t_{ij}^{\mathcal{X}}$ and $t_{ji}^{\mathcal{Y}}$ are continuously differentiable and thus Newton-differentiable, as well as strictly differentiable. By applying the chain rule twice, we obtain Newton- or strict differentiability of $\psi_{j}\circ F\circ\phi_{j}^{-1}$. ∎
Using independence of chart representations, we can finally define Newton-differentiability for mappings between manifolds:
@@
We call $F$ Newton-differentiable at a point $x_{0}\in\mathcal{X}$ with respect to $F^{\prime}$ if all chart representations of $F$ and the corresponding representations of $F^{\prime}$ satisfy (7) for $\xi_{0}=\phi(x_{0})$. In this case we call $F^{\prime}$ a Newton-derivative of $F$ at $x_{0}$.
Analogously, we can define strict differentiability:
@@
Let $F:\mathcal{X}\to\mathcal{Y}$. $F$ is called strictly differentiable at $x_{0}\in\mathcal{X}$, if its derivative $F^{\prime}(x_{0})\in L(T_{x_{0}}\mathcal{X},T_{F(x_{0})}\mathcal{Y})$ satisfies (8) in all chart representations for $\xi_{0}=\phi(x_{0})$.